Complete the inverse of F(x) below. Enter exponents using the caret (^); for example, enter x2 as x^2. Do not include "F-1(y) =" in your answer.
F(x) = x - 4
F-1(y) = _____

can some one explan to me how to do it step by step for further refrence

Answer by nabla(475)

write it out as:
y=x-4 first
now switch the x and y
x=y-4
now solve for y
y=x+4

x+4 would be the part you enter as the answer.